A video game cost $49 when it first went on sale. After three months, the price of the game changed by k percent, where a positive value of k indicates a price increase. If the new price of the game was $35, which equation correctly relates the old and new prices of the game?

Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

the cost of the video game was 49 dollars when it first went on sale.
after 3 months the price of the game changed by k percent.
if the new price of the game was 35 dollars, which equation correctly relates the old and new prices of the game?

to solve for k, you would use the equation 49 - k = 35
add k to both sides of this equation and subtract 35 from both sides of this equation to get 49 - 35 = k
solve for k to get k = 14 dollars.

if the equation was 49 + k = 35, then you would do the same thing, except now that you should get a negative value for k.
subtract k from both sides of the equation and subtract 35 from both sides of the equation to get 49 - 35 = -k
switch sides to get -k = 49 - 35
simplify to get -k = 14
solve for k by multiplying both sides of the equation by -1 to get k = -14.

since i don't know what your options are, i have to go with what i just showed you.

your equation could have been 49 - k = 35 or it could have been 49 + k = 35.
49 - k leads to a positive value for k and 49 + k leads to a negative value for k.
both are valid, depending on whether you want the value of k to be positive or negative.
if you want the value of k to be posiive, then you have to go with 49 - k.

if you send me the equations that you are dealing with, i can probably pick the correct one out for you.

theo

Answer by MathTherapy(10552)


A video game cost $49 when it first went on sale. After three months, the price of the game changed by k percent, where a positive value of k indicates a price increase. If the new price of the game was $35, which equation correctly relates the old and new prices of the game?
"which equation correctly relates the old and new prices of the game?" almost certainly should include a list of choices.
Don't you think you need to provide this?